# Index Terms:

### v1

@@ Appendix C Proof of Theorem 1
where the expectation is over $\bm{r}$; the randomness is introduced from $\nu$ and ${\bm{A}}$. We can write
@@
Consider (b) first. Note that $W_{i}=-\sum_{q\in\mathcal{U}}\mathbb{1}(r_{i}=q)\log(f_{q}(x_{i}^{\natural}))$ for all $i$ are independent, where $W_{i}\in[0,U_{f}]$ holds. Then, for any $10<\eta<$ 1, with probability at least $1-\eta$, we have:
@@
Term (a) amounts to the difference between sample average ${\cal L}({\bm{x}}^{\star})$ and the expected value $\mathbb{E}[{\cal L}({\bm{x}}^{\star})]$. Hence, we can bound term (a) by the Rademacher complexity bounds [35, Theorem 26.5]. With probability of at least $1-\delta$, we have
@@
^$\widehat{\mathcal{R}}(\mathcal{W}):=\mathbb{E}_{\bm{\sigma}}\left[\underset{w\in\mathcal{W}}{\sup}\frac{1}{R}\sum_{i=1}^{R}\sigma_{i}w(r_{i})\right]$ (42)
@@
^$\widehat{\mathcal{R}}({\cal W})\leq L_{f}\widehat{\mathcal{R}}(\mathcal{S}),$ (43)
@@
^$\widehat{\mathcal{R}}(\mathcal{S})\leq\xi(\mathcal{S}),$ (45)
[…]
@@
From [55, Corollary 5.35], we have
@@
with probability at least $1-2e^{-R/2}$. We can choose $\mu=2+\sqrt{\nicefrac{{J}}{{R}}}$. Then, we arrive at $\xi(\mathcal{S})\leq\frac{4\mu}{\sqrt{R}}+\frac{12}{\sqrt{R}}\sqrt{{4K}\log\left(2\sqrt{2\pi\kappa}L_{\cal G}\right)}.$
This result, combined with (41), (43) and (45), yields $\mathbb{E}[{\cal L}({\bm{x}}^{\star})]-{\cal L}({\bm{x}}^{\star})\leq 2\frac{L_{f}}{\sqrt{R}}\tau+4U_{f}\sqrt{\frac{2\log(4/\delta)}{R}}$ with probability at least $1-\delta-2e^{-R/2}$, where $\tau=4\left(2+\sqrt{\nicefrac{{J}}{{R}}}\right)+12\sqrt{{4K}\log\left(2\sqrt{2\pi\kappa}L_{\cal G}\right)}.$
@@
$\begin{split}&{\rm KL}({\bm{x}}^{\natural}||{\bm{x}}^{\star})\\ \leq&2\frac{L_{f}}{\sqrt{R}}\tau+4U_{f}\sqrt{\frac{2\log(2/\eta)}{R}}+\frac{3}{2}\nu\frac{L_{f}}{\sqrt{R}}+U_{f}\sqrt{\frac{1}{2R}\log\frac{2}{\eta}}\\ =&2\frac{L_{f}}{\sqrt{R}}(\tau+\frac{3}{4}\nu)+\frac{9}{2}\frac{U_{f}}{\sqrt{R}}\sqrt{\log(2/\eta)}\\ =&\frac{1}{\sqrt{R}}\left(2L_{f}(\tau+\frac{3}{4}\nu)+\frac{9}{2}U_{f}\sqrt{\log(2/\eta)}\right)\end{split}$
with probability at least $1-3\eta-e^{-\Omega(R)}$.
@@
with $10<\gamma<$ 1 and $\epsilon>0$, with probability $1-e^{-\Omega((1-\gamma)^{2}R)}$.
@@
The proof is in the same vein of [45, Lemma 4.1]. The difference is as follows: [45, Lemma 4.1] is for the case where the parameter space is an $\ell_{2}$-norm ball, while in our case, the parameter space is the bounded manifold $\Gamma$, which is compact. We replace the converting number for the $\ell_{2}$-norm ball in [45, Lemma 4.1] by the covering number for the manifold $\Gamma$, and the remaining proof remains the same. This leads to the desired result in Lemma 1. ∎
@@
By substituting (47) into (49) and using $R\leq 4MN$, one can obtain $\nicefrac{{\|{\bm{A}}({\bm{h}}^{\natural}-{\bm{h}}^{\star})\|_{2}^{2}}}{{MN}}\leq\nicefrac{{8}}{{\sqrt{R}}}(4\nicefrac{{L_{f}}}{{F_{f}}}(\tau+\nicefrac{{3}}{{4}}\nu)+9\nicefrac{{U_{f}}}{{F_{f}}}\sqrt{\log(2/\eta)}),$ and $\nicefrac{{\epsilon^{\prime 2}}}{{MN}}\leq\nicefrac{{4\epsilon^{\prime 2}}}{{R}}$. Choosing $\gamma=1/2$ and $\epsilon=1/K$, we obtain the result in Theorem 1, with probability at least $1-3\eta-e^{-\Omega(R)}$.
@@
Appendix D Proof of Fact 1
To prove that the mapping $\cal G$ is Lipschitz continuous, it suffices to show the Jacobian of $\cal G$ is bounded; see, e.g., [57, Theorem 9.7]. Let us denote the Jacobian matrix w.r.t. $\theta_{k}$ as $\nabla_{\theta_{k}}\bm{H}\in\mathbb{R}^{M\times N}$, where $[\nabla_{\theta_{k}}\bm{H}]_{m,n}=\nicefrac{{\partial[\bm{H}]_{m,n}}}{{\partial\theta_{k}}}$. Then, it is readily seen that
@@ Appendix D Proof of Fact 1
Appendix E Proof of Theorem 2
@@
(S-REC of ${\bm{D}}$) Let the compressing matrix ${\bm{D}}\in\mathbb{R}^{N_{S}\times J}$ and
${\bm{D}}=\frac{1}{N_{S}}({\bm{S}}^{\top}\odot{\bm{A}}^{\top})^{\top},$
@@
$\begin{split}&s_{i,j}\sim{\cal N}(0,1),\quad i=1,\ldots,2N,\quad j=1,\ldots,N_{S},\\ &a_{i,j}\sim{\cal N}(0,1),\quad i=1,\ldots,2M,\quad j=1,\ldots,N_{S}\end{split}$
@@
$N_{S}=\Omega\left(K^{2}\left(\log\frac{4\sqrt{2}\sqrt{\pi\kappa}L_{\cal G}}{\epsilon}\right)^{2}\right).$
@@
with $10<\gamma<$ 1 and $\epsilon>0$, with probability $1-e^{-\Omega((1-\gamma)^{2}N_{S})}$.
The proof for Lemma 2 is shown in Section F. With Lemma 2, the remaining proof is the same as those for Theorem 1.
Appendix F Proof of Lemma 2
@@
Also, for any $\epsilon>0$ and fixed $N_{S}$, if
$\alpha\geq\max\left(\sqrt{\frac{c_{1}}{N_{S}}\log\frac{1}{\epsilon}},\frac{c_{2}^{2}}{N_{s}}\left(\log\frac{1}{\epsilon}\right)^{2}\right),$
@@
Let the compressing matrix ${\bm{D}}\in\mathbb{R}^{N_{S}\times J}$ and
${\bm{D}}=\frac{1}{N_{S}}({\bm{S}}^{\top}\odot{\bm{A}}^{\top})^{\top},$
@@
$\begin{split}&s_{i,j}\sim{\cal N}(0,1),\quad i=1,\ldots,2N,\quad j=1,\ldots,N_{S},\\ &a_{i,j}\sim{\cal N}(0,1),\quad i=1,\ldots,2M,\quad j=1,\ldots,N_{S}\end{split}$
@@
$N_{S}=\Omega\left(K^{2}\left(\log\frac{4\sqrt{2}\sqrt{\pi\kappa}L_{\cal G}}{\epsilon}\right)^{2}\right).$
@@
then we have $\|{\bm{D}}(\check{{\bm{h}}}-\check{{\bm{h}}}^{\prime})\|_{2}={O}(\epsilon)$ with probability at least $1-e^{-\Omega(\sqrt{N_{S}})}$.
The proof of Lemma 4 will be shown later. Now we are ready to show the main proof of Lemma 2. First, we define $\check{s}=\check{g}\circ{\cal G}$. We construct an $\epsilon/2L_{\cal G}$-net $M$ of $\mathcal{Z}$. Then $N=\check{s}(M)$ is an $\epsilon$-net of $\check{\Gamma}$. Let $T$ denote the set of pairwise distance of $N$, i.e., $T=\{\check{s}({\bm{z}}_{1})-\check{s}({\bm{z}}_{2})\mid{\bm{z}}_{1},{\bm{z}}_{2}\in M\}.$
Then, the cardinality $|T|=|N|^{2}$ and
$\log(|T|)\leq 8K\log\left(\frac{4\sqrt{2}\sqrt{\pi\kappa}L_{\cal G}}{\epsilon}\right).$ (50)
@@
with probability of at least $1-e^{-{\mathbf{\Omega}(\sqrt{N_{s}})}}$.
@@
Appendix G Proof of Lemma 3
@@
Following the result in [58, Theorem 42] with the special case $c=2$, it can be verified that the ${\bm{D}}$ satisfies the $(\alpha,\epsilon,p)$-JL-moment property for any $p\geq 4$ if
$N_{s}\geq\max(c_{1}\alpha^{-2}\log 1/\epsilon,c_{2}\alpha^{-1}(\log 1/\epsilon)^{2})$
@@
Appendix H Proof of Lemma 4
From Lemma 3, we have $\Pr(\|{\bm{D}}{\bm{x}}\|_{2}\geq(1+\alpha)\|{\bm{x}}\|_{2})\leq f$ for any $f>0$ if $\alpha\geq\max\big{\{}\sqrt{\frac{c_{1}}{N_{S}}\log(1/f)},\frac{c_{2}^{2}}{N_{S}}(\log(1/f))^{2}\big{\}}$.
@@ Appendix H Proof of Lemma 4
[…]
@@
$T_{i}=\{\check{{\bm{h}}}_{i+1}-\check{{\bm{h}}}_{i}|\check{{\bm{h}}}_{i+1}\in N_{i+1},\check{{\bm{h}}}_{i}\in N_{i}\}$
Then, the cardinality of each $T_{i}$ can be bounded as follows
$\begin{split}\log|T_{i}|\leq&\log|N_{i+1}|+\log|N_{i}|\\ \leq&12iK+8K\log\left(\frac{4\sqrt{2}\sqrt{\pi\kappa}L_{\cal G}}{\epsilon}\right)\end{split}$
@@
$\Pr(\|{\bm{D}}\check{{\bm{h}}}\|_{2}\leq(1+\alpha_{i})\|\check{{\bm{h}}}\|_{2},\forall\check{{\bm{h}}}\in T_{i})\geq 1-\sum_{i=0}^{l-1}|T_{i}|f_{i}$
@@
Now, let $N_{s}^{1/2}=12K\log\left(\frac{4\sqrt{2}\sqrt{\pi\kappa}L_{\cal G}}{\epsilon}\right),$ and $\log f_{i}=-(N_{s}^{1/2}+16iK),$ which leads to
$\alpha_{i}$ $=\max\left(\frac{c_{1}^{1/2}}{N_{s}^{1/2}}(N_{s}^{1/2}+16iK)^{1/2},\frac{c_{2}^{2}}{N_{s}}(N_{s}^{1/2}+16iK)^{2}\right)$
@@
$\log(|T_{i}|f_{i})$ $\leq 12iK+\frac{2}{3}N_{s}^{1/2}-N_{s}^{1/2}-16iK$
$=-4iK-\frac{\sqrt{N_{s}}}{3}$
@@
$\Pr({\bm{D}}\check{{\bm{h}}}_{2}\leq(1+\alpha_{i})\|\check{{\bm{h}}}\|_{2},\forall\check{{\bm{h}}}\in T_{i})$
$\geq$ $1-\sum_{i=0}^{l-1}e^{-4iK-\sqrt{N_{s}}/3}\geq 1-e^{-\sqrt{N_{s}}/3}\left(\frac{1}{1-e^{-1}}\right)$
$\geq$ $1-2e^{-\sqrt{N_{s}}/3}$
@@
We have ${\bm{h}}_{i+1}-{\bm{h}}_{i}\in T_{i}$, and with probability of at least $1-2e^{-\sqrt{N_{s}}/3}$, the following holds
@@
Since $\alpha_{i}$ decreases with value of $N_{s}$ and the term inside summation is decreasing, it can be upper bounded as
@@
Then, for any $\check{{\bm{h}}},\check{{\bm{h}}}^{\prime}\in\check{\Gamma}$, with probability of at least $1-e^{-\mathbf{\Omega}(\sqrt{N_{s}})}$, we have
@@
Let ${\bm{D}}=({\bm{A}}_{1}^{\top}\odot{\bm{A}}_{2}^{\top})^{\top},$ where $\odot$ denotes the Khatri-Rao product, $[{\bm{A}}_{1}]_{i,j}\sim{\cal N}(0,1/N_{S})$ and $[{\bm{A}}_{2}]_{i,j}\sim{\cal N}(0,1/N_{S})$. Then, with probability at least $1-e^{-N_{A}/2}$, it holds that
$\|{\bm{D}}\|_{2}\leq 4+4\sqrt{\frac{J}{N_{S}}}+\frac{J}{N_{S}}.$ (56)
The proof of Lemma 5 will be shown soon. Using result from Lemma 5, with probability of at least $1-4e^{-N_{s}/2}$, we have $\|{\bm{D}}\check{{\bm{h}}}^{f}\|_{2}\leq\left(4+4\sqrt{\frac{{J}}{N_{s}}}+\frac{{J}}{N_{s}}\right)\frac{\epsilon}{2^{l}}.$
@@
with probability of at least $1-e^{-{\mathbf{\Omega}(\sqrt{N_{s}})}}$.
Appendix I Proof of Lemma 5
@@
From [55, Corollary 5.35], for both ${\bm{A}}_{1}$ and ${\bm{A}}_{2}$, we have $\|{\bm{A}}_{i}\|_{2}\geq 2+\sqrt{n_{i}/N_{s}}$ with probability at most $2e^{-N_{s}/2}$ where, $n_{1}=2N$ and $n_{2}=2M$. Now, by using union bound we have

### v2

@@ VII Conclusion
There are several promising directions for future exploration. First, one could investigate more signal-adaptive designs for the compression matrix. In this work, we used random compression matrices primarily to preserve properties from the JL Lemma, which was central to our recoverability analysis. However, leveraging structure in the channel or data to design tailored compression matrices could further reduce feedback overhead—an appealing and practical direction. Second, on the theoretical front, more fine-grained characterizations could be developed—for instance, clarifying the relationships between $U_{f}$, $L_{f}$, and $F_{f}$, and how they depend on the quantization level. Third, from an implementation standpoint, integrating online optimization or learning-based techniques could significantly enhance computational efficiency and scalability.
@@ Appendix C Proof of Theorem 1
where the expectation is over $\bm{r}$; the randomness is introduced from $\bm{v}$} and $\bm{A}$}. We can write
@@
Consider (b) first. Note that $W_{i}=-\sum_{q\in\mathcal{U}}\mathbb{1}(r_{i}=q)\log(f_{q}(x_{i}^{\natural}))$ for all $i$ are independent, where $W_{i}\in[0,U_{f}]$ holds. Then, for any $0<\eta<1$, with probability at least $1-\eta{-\vartheta}$, we have:
@@
Term (a) amounts to the difference between sample average ${\cal L}({\bm{x}}^{\star})$ and the expected value $\mathbb{E}[{\cal L}({\bm{x}}^{\star})]$. In addition, note that all elements of $\bm{x}$} is bounded with probability at least $1-\vartheta$. Hence, we can bound term (a) by the Rademacher complexity bounds [35, Theorem 26.5]. With probability of at least $1-\delta{-\vartheta}$, we have
@@
$\widehat{\mathcal{R}}(\mathcal{W}):=\mathbb{E}_{\bm{\sigma}}\left[\underset{w\in\mathcal{W}}{\sup}\frac{1}{R}\sum_{i=1}^{R}\sigma_{i}w(r_{i})\right]$ (44)
@@
$\widehat{\mathcal{R}}({\cal W})\leq L_{f}\widehat{\mathcal{R}}(\mathcal{S}),$ (45)
@@
$\widehat{\mathcal{R}}(\mathcal{S})\leq\xi(\mathcal{S}),$ (47)
where $\xi(\mathcal{S})=\inf_{\mu>0}\left(\frac{4\mu}{\sqrt{R}}+\frac{12}{R}\int_{\mu}^{\sqrt{R}}\sqrt{\log\mathcal{C}(\mathcal{S},\epsilon)}\ d\epsilon\right).$
Since $\sqrt{\log\mathcal{C}(\mathcal{S},\mu)}$ increases as $\mu$ decreases, the integral can be upper bounded by $\int_{\mu}^{\sqrt{R}}\sqrt{\log\mathcal{C}(\mathcal{S},\epsilon)}\ d\epsilon\leq\sqrt{R}\sqrt{\log\mathcal{C}(\mathcal{S},\mu)}.$ As a result, we can bound $\xi(\mathcal{S})\leq\inf_{\mu>0}\left(\frac{4\mu}{\sqrt{R}}+\frac{12}{R}\sqrt{R}\sqrt{\log\mathcal{C}(\mathcal{S},\mu)}\right).$
@@
From [57, Corollary 5.35], we have
@@
with probability at least $1-2e^{-R/2}$. We can choose $\mu=2+\sqrt{\nicefrac{{J}}{{R}}}$. Then, we arrive at $\xi(\mathcal{S})\leq\frac{4{(2+\sqrt{J/R})}}{\sqrt{R}}+\frac{12}{\sqrt{R}}\sqrt{{4K}\log\left(2\sqrt{2\pi\kappa}L_{\cal G}\right)}.$
This result, combined with (43), (45) and (47), yields $\mathbb{E}[{\cal L}({\bm{x}}^{\star})]-{\cal L}({\bm{x}}^{\star})\leq 2\frac{L_{f}}{\sqrt{R}}\tau+4U_{f}\sqrt{\frac{2\log(4/\delta)}{R}}$ with probability at least $1-\delta{-\vartheta}-2e^{-R/2}$, where $\tau=4\left(2+\sqrt{\nicefrac{{J}}{{R}}}\right)+12\sqrt{{4K}\log\left(2\sqrt{2\pi\kappa}L_{\cal G}\right)}.$
@@
$\begin{split}&{\rm KL}({\bm{x}}^{\natural}||{\bm{x}}^{\star})\\ \leq&2\frac{L_{f}}{\sqrt{R}}\tau+4U_{f}\sqrt{\frac{2\log(2/\eta)}{R}}+\frac{3}{2}\nu\frac{L_{f}}{\sqrt{R}}+U_{f}\sqrt{\frac{1}{2R}\log\frac{2}{\eta}}\\ =&2\frac{L_{f}}{\sqrt{R}}(\tau+\frac{3}{4}\nu)+\frac{9}{{\sqrt{2}}}\frac{U_{f}}{\sqrt{R}}\sqrt{\log(2/\eta)}\\ =&\frac{1}{\sqrt{R}}\left(2L_{f}(\tau+\frac{3}{4}\nu)+\frac{9}{{\sqrt{2}}}U_{f}\sqrt{\log(2/\eta)}\right)\end{split}$
with probability at least $1-3\eta{-\vartheta}-e^{-\Omega(R)}$.
@@
with $0<\gamma<1$ and $\epsilon>0$, with probability $1-e^{-\Omega((1-\gamma)^{2}R)}$.
@@
The proof is in the same vein of [45, Lemma 4.1]. The difference is as follows: [45, Lemma 4.1] is for the case where the parameter space is an $\ell_{2}$-norm ball, while in our case, the parameter space is the bounded manifold $\Gamma$, which is compact. We replace the covering number for the $\ell_{2}$-norm ball in [45, Lemma 4.1] by the covering number for the manifold $\Gamma$, and the remaining proof remains the same. This leads to the desired result in Lemma 1. ∎
@@
By substituting (49) into (51) and using $R\leq 4MN$, one can obtain $\nicefrac{{\|{\bm{A}}({\bm{h}}^{\natural}-{\bm{h}}^{\star})\|_{2}^{2}}}{{MN}}\leq\nicefrac{{8}}{{\sqrt{R}}}(4\nicefrac{{L_{f}}}{{F_{f}}}(\tau+\nicefrac{{3}}{{4}}\nu)+9\nicefrac{{U_{f}}}{{F_{f}}}\sqrt{{2}\log(2/\eta)}),$ and $\nicefrac{{\epsilon^{\prime 2}}}{{MN}}\leq\nicefrac{{4\epsilon^{\prime 2}}}{{R}}$. Choosing $\gamma=1/2$ and $\epsilon=1/K$, we obtain the result in Theorem 1, with probability at least $1-3\eta{-\vartheta}-e^{-\Omega(R)}$.
@@
Appendix D Per-iteration complexity of REDEEM at BS
The per-iteration complexity of our proposed Scheme 1 algorithm is analyzed below:
  1. 1.
    Dual variable update: This step involves only matrix additions. The time complexity is $O(MN)$.
  2. 2.
    EM update: Note that the expensive inverse operation in (37) doesn’t depend on the optimization variables, and it can be just computed once in the beginning. The time complexities of E step and the M step are $O(RMN)$ and $O(MNR)$, respectively.
  3. 3.
    Mod-RELAX update: This step uses 2D-DFT for solving (33) whose performance is reliant on the discretization resolution of the DFT matrix. Our implementation has a time complexity of $O(\bar{I}_{GA}\bar{I}_{RLX}\bar{I}_{RLX}K^{3}MN+\bar{I}_{RLX}K^{2}MN\log MN)$.
  4. 4.
    Total per-iteration complexity is $O(\bar{I}_{\text{EM}}MNR+\bar{I}_{GA}\bar{I}_{RLX}K^{3}MN+\bar{I}_{RLX}K^{2}MN\log MN)$.
In the above, $\bar{I}_{\text{GA}},\bar{I}_{\text{EM}}$ and $\bar{I}_{\text{RLX}}$ denote the maximum number of iterations for gradient ascent, EM, and Mod-RELAX, respectively. Empirically, the algorithm converges in a few iterations. Note that Scheme 2 has the same per-iteration complexity as that of Scheme 1, with the measurement dimension $R$ simply replaced by $T$.
Appendix E Proof of Fact 1
To prove that the mapping $\cal G$ is Lipschitz continuous, it suffices to show the Jacobian of $\cal G$ is bounded; see, e.g., [59, Theorem 9.7]. Let us denote the Jacobian matrix w.r.t. $\theta_{k}$ as $\nabla_{\theta_{k}}\bm{H}\in\mathbb{R}^{M\times N}$, where $\nabla_{\theta_{k}}\bm{H}]_{m,n}=\nicefrac{{\partial[\bm{H}]_{m,n}}}{{\partial\theta_{k}}}$. Then, it is readily seen that
@@ Appendix E Proof of Fact 1
Appendix F Proof of Theorem 2
@@
(S-REC of $\bm{D}$}) Let the compressing matrix ${\bm{D}}\in\mathbb{R}^{{T}\times J}$ and
${\bm{D}}=\frac{1}{{T}}({\bm{S}}^{\top}\odot{\bm{A}}^{\top})^{\top},$
@@
$\begin{split}&s_{i,j}\sim{\cal N}(0,1),\quad i=1,\ldots,2N,\quad j=1,\ldots,{T},\\ &a_{i,j}\sim{\cal N}(0,1),\quad i=1,\ldots,2M,\quad j=1,\ldots,{T}\end{split}$
@@
${T}=\Omega\left(K^{2}\left(\log\frac{4\sqrt{2}\sqrt{\pi\kappa}L_{\cal G}}{\epsilon}\right)^{2}\right).$
@@
with $0<\gamma<1$ and $\epsilon>0$, with probability $1-e^{-\Omega((1-\gamma)^{2}{T})}$.
The proof for Lemma 2 is shown in Section G. With Lemma 2, the remaining proof is the same as those for Theorem 1.
Appendix G Proof of Lemma 2
@@
Also, for any $\epsilon>0$ and fixed T{T}, if
$\alpha\geq\max\left(\sqrt{\frac{c_{1}}{{T}}\log\frac{1}{\epsilon}},\frac{c_{2}^{2}}{{T}}\left(\log\frac{1}{\epsilon}\right)^{2}\right),$
@@
Let the compressing matrix ${\bm{D}}\in\mathbb{R}^{{T}\times J}$ and
${\bm{D}}=\frac{1}{{T}}({\bm{S}}^{\top}\odot{\bm{A}}^{\top})^{\top},$
@@
$\begin{split}&s_{i,j}\sim{\cal N}(0,1),\quad i=1,\ldots,2N,\quad j=1,\ldots,{T},\\ &a_{i,j}\sim{\cal N}(0,1),\quad i=1,\ldots,2M,\quad j=1,\ldots,{T}\end{split}$
@@
${T}=\Omega\left(K^{2}\left(\log\frac{4\sqrt{2}\sqrt{\pi\kappa}L_{\cal G}}{\epsilon}\right)^{2}\right).$
@@
then we have $\|{\bm{D}}(\check{{\bm{h}}}-\check{{\bm{h}}}^{\prime})\|_{2}={O}(\epsilon)$ with probability at least $1-e^{-\Omega(\sqrt{{T}})}$.
The proof of Lemma 4 will be shown later. Now we are ready to show the main proof of Lemma 2. First, we define $\check{s}=\check{g}\circ{\cal G}$. We construct an $\epsilon/2L_{\cal G}$-net $M$ of $\mathcal{Z}$. Then $N=\check{s}(M)$ is an $\epsilon$-net of $\check{\Gamma}$. Let $U$ denote the set of pairwise distance of $N$, i.e., $U=\{\check{s}({\bm{z}}_{1})-\check{s}({\bm{z}}_{2})\mid{\bm{z}}_{1},{\bm{z}}_{2}\in M\}.$
Then, the cardinality $|U|=|N|^{2}$ and
$\log(|U|)\leq 8K\log\left(\frac{4\sqrt{2}\sqrt{\pi\kappa}L_{\cal G}}{\epsilon}\right).$ (52)
@@
with probability of at least $1-e^{-{\mathbf{\Omega}(\sqrt{{T}})}}$.
@@
Appendix H Proof of Lemma 3
@@
Following the result in [60, Theorem 42] with the special case $c=2$, it can be verified that the $\bm{D}$} satisfies the $(\alpha,\epsilon,p)$-JL-moment property for any $p\geq 4$ if
${T}\geq\max(c_{1}\alpha^{-2}\log 1/\epsilon,c_{2}\alpha^{-1}(\log 1/\epsilon)^{2})$
@@
Appendix I Proof of Lemma 4
From Lemma 3, we have $\Pr(\|{\bm{D}}{\bm{x}}\|_{2}\geq(1+\alpha)\|{\bm{x}}\|_{2})\leq f$ for any $f>0$ if $\alpha\geq\max\big{\{}\sqrt{\frac{c_{1}}{{T}}\log(1/f)},\frac{c_{2}^{2}}{{T}}(\log(1/f))^{2}\big{\}}$.
@@ Appendix I Proof of Lemma 4
$\begin{split}\log|M_{i}|\leq&\leavevmode\nobreak\ 4K\log\left(\frac{4\sqrt{2}\sqrt{\pi\kappa}L_{\cal G}}{\epsilon_{i}}\right)\\ \leq&\leavevmode\nobreak\ 4iK+4K\log\left(\frac{4\sqrt{2}\sqrt{\pi\kappa}L_{\cal G}}{\epsilon}\right).\end{split}$
@@
${U_{i}}=\{\check{{\bm{h}}}_{i+1}-\check{{\bm{h}}}_{i}|\check{{\bm{h}}}_{i+1}\in N_{i+1},\check{{\bm{h}}}_{i}\in N_{i}\}$
Then, the cardinality of each $U_{i}$ can be bounded as follows
$\begin{split}{\log|U_{i}|}\leq&\log|N_{i+1}|+\log|N_{i}|\\ \leq&12iK+8K\log\left(\frac{4\sqrt{2}\sqrt{\pi\kappa}L_{\cal G}}{\epsilon}\right)\end{split}$
@@
$\Pr(\|{\bm{D}}\check{{\bm{h}}}\|_{2}\leq(1+\alpha_{i})\|\check{{\bm{h}}}\|_{2},\forall\check{{\bm{h}}}\in{U_{i}})\geq 1-\sum_{i=0}^{l-1}|{U_{i}}|f_{i}$
@@
Now, let ${T}^{1/2}=12K\log\left(\frac{4\sqrt{2}\sqrt{\pi\kappa}L_{\cal G}}{\epsilon}\right),$ and $\log f_{i}=-({T}^{1/2}+16iK),$ which leads to
$\alpha_{i}$ $=\max\left(\frac{c_{1}^{1/2}}{{T}^{1/2}}({T}^{1/2}+16iK)^{1/2},\frac{c_{2}^{2}}{{T}}({T}^{1/2}+16iK)^{2}\right)$
@@
$\log(|{U_{i}}|f_{i})$ $\leq 12iK+\frac{2}{3}{T}^{1/2}-{T}^{1/2}-16iK$
$=-4iK-\frac{\sqrt{{T}}}{3}$
@@
$\Pr({\bm{D}}\check{{\bm{h}}}_{2}\leq(1+\alpha_{i})\|\check{{\bm{h}}}\|_{2},\forall\check{{\bm{h}}}\in{U_{i}})$
$\geq$ $1-\sum_{i=0}^{l-1}e^{-4iK-\sqrt{{T}}/3}\geq 1-e^{-\sqrt{{T}}/3}\left(\frac{1}{1-e^{-1}}\right)$
$\geq$ $1-2e^{-\sqrt{{T}}/3}$
@@
We have ${\bm{h}}_{i+1}-{\bm{h}}_{i}\in{U_{i}}$, and with probability of at least $1-2e^{-\sqrt{{T}}/3}$, the following holds
@@
Since $\alpha_{i}$ decreases with value of T{T} and the term inside summation is decreasing, it can be upper bounded as
@@
Then, for any $\check{{\bm{h}}},\check{{\bm{h}}}^{\prime}\in\check{\Gamma}$, with probability of at least $1-e^{-\mathbf{\Omega}(\sqrt{{T}})}$, we have
@@
Let ${\bm{D}}=({\bm{A}}_{1}^{\top}\odot{\bm{A}}_{2}^{\top})^{\top},$ where $\odot$ denotes the Khatri-Rao product, ${\bm{A}}_{1}]_{i,j}\sim{\cal N}(0,1/{T})$ and ${\bm{A}}_{2}]_{i,j}\sim{\cal N}(0,1/{T})$. Then, with probability at least $1-4e^{-{T}/2}$, it holds that
$\|{\bm{D}}\|_{2}\leq 4+4\sqrt{\frac{J}{{T}}}+\frac{J}{{T}}.$ (58)
The proof of Lemma 5 will be shown soon. Using result from Lemma 5, with probability of at least $1-4e^{-{T}/2}$, we have $\|{\bm{D}}\check{{\bm{h}}}^{f}\|_{2}\leq\left(4+4\sqrt{\frac{{J}}{{T}}}+\frac{{J}}{{T}}\right)\frac{\epsilon}{2^{l}}.$
@@
with probability of at least $1-e^{-{\mathbf{\Omega}(\sqrt{{T}})}}$.
Appendix J Proof of Lemma 5
@@
From [57, Corollary 5.35], for both ${\bm{A}}_{1}$ and ${\bm{A}}_{2}$, we have $\|{\bm{A}}_{i}\|_{2}\geq 2+\sqrt{n_{i}/{T}}$ with probability at most $2e^{-{T}/2}$ where, $n_{1}=2N$ and $n_{2}=2M$. Now, by using union bound we have